### v2

# The metric compactification of a Kobayashi hyperbolic complex manifold and a Denjoy–Wolff Theorem

Abstract.
We study the metric compactification of a Kobayashi hyperbolic complex manifold $\mathcal{X}$ equipped with the Kobayashi distance $\mathsf{k}_{\mathcal{X}}$. We show that this compactification is genuine — i.e., $\mathcal{X}$ embeds as a dense open subset — even without completeness of $\mathsf{k}_{\mathcal{X}}$, and that it becomes a good compactification in the sense of Bharali–Zimmer when $(\mathcal{X},\mathsf{k}_{\mathcal{X}})$ is complete. As an application, we obtain a criterion for the continuous extension of quasi-isometric embeddings from $(\mathcal{X},\mathsf{k}_{\mathcal{X}})$ into visibility domains of complex manifolds. For a Kobayashi hyperbolic domain $\Omega\subsetneq\mathcal{X}$, to each boundary point of $\Omega$ in the end compactification, we associate a fiber of metric boundary points. This allows the small and big horospheres of Abate to be expressed as the intersection and union of horoballs centered at metric boundary points. We use this to formulate a Wolff-type lemma in terms of horoballs and prove a Denjoy–Wolff theorem for complete hyperbolic domains satisfying a boundary divergence condition for $\mathsf{k}_{\mathcal{X}}$. Finally, we present necessary and sufficient conditions under which the identity map extends continuously between the metric and end compactifications.
Key words and phrases:
Kobayashi hyperbolic domain, metric compactification, end compactification, geodesic visibility, Denjoy–Wolff theorem
2020 Mathematics Subject Classification:
Primary: 32F45, 30D40; Secondary: 32Q45, 53C23.
1. Introduction: Statement of main results
[…]
Definition 1.1.
Let $(X,d)$ be a metric space. A pair $(\iota,X^{*})$ is said to be a compactification of $X$ if $X^{*}$ is a sequentially compact Hausdorff topological space, $\iota:X\longrightarrow X^{*}$ is a homeomorphism onto its image, and $\iota(X)$ is open and dense in $X^{*}$.
[…]
Theorem 1.2.
Let $\mathcal{X}$ be a Kobayashi hyperbolic complex manifold equipped with the Kobayashi distance $\mathsf{k}_{\mathcal{X}}$. Then $\big{(}\overline{\mathcal{X}}^{\mathsf{k}},i_{\mathsf{k}}\big{)}$ is a compactification of $\mathcal{X}$.
Here, $\overline{\mathcal{X}}^{\mathsf{k}}$ denotes the metric compactification as defined in Subsection 2.2, and $i_{\mathsf{k}}:\mathcal{X}\to\overline{\mathcal{X}}^{\mathsf{k}}$ is the canonical embedding. We emphasize that no completeness assumption is imposed on $\mathsf{k}_{\mathcal{X}}$. Indeed, for general hyperbolic complex manifolds, the completeness of the Kobayashi distance is often difficult to verify. In the absence of geodesics, our proof instead relies on the existence of $(1,\epsilon)$-quasi-geodesics connecting any two points in $\mathcal{X}$ for arbitrary $\epsilon>0$ — a fact established by Masanta [17, Proposition 2.8], adapting the argument used by Bharali–Zimmer [6, Proposition 4.4] for bounded domains.
In the same paper, Bharali–Zimmer also introduced the notion of a good compactification (Definition 3.3) for geodesic metric spaces that captures the idea that geodesics converging to the same ideal boundary point must eventually leave every compact subset. Such a condition rules out geodesic looping near the boundary and plays a key role in the context of continuous extension of quasi-isometric embeddings; [6, Section 6.2]. Using results of Webster–Winchester [20], we observe the following:
Proposition 1.3.
[…]
The Hopf–Rinow theorem implies that a complete Kobayashi hyperbolic complex manifold $\mathcal{X}$ is proper and geodesic, and hence the above proposition applies to it. For a domain $D\Subset\mathbb{C}^{n}$ such that $(D,\mathsf{k}_{D})$ admits a good compactification $(D^{*},\iota)$, Bharali–Zimmer [6, Theorem 6.5] showed that any quasi-isometric embedding $F:D\longrightarrow\Omega$ — where $\Omega$ is a visibility domain — extends continuously to a map from $D^{*}$ to $\overline{\Omega}$. As an application of Proposition 1.3, we obtain:
Theorem 1.4.
Let $\mathcal{X}$ be a complete Kobayashi hyperbolic complex manifold, and let $\Omega\subsetneq\mathcal{Y}$ be a Kobayashi hyperbolic domain in a complex manifold $\mathcal{Y}$ that satisfies the visibility property. Then any continuous quasi-isometric embedding $F:\mathcal{X}\longrightarrow\Omega$ extends to a continuous map from $\overline{\mathcal{X}}^{\mathsf{k}}$ to $\overline{\Omega}^{\rm End}$.
Here, visibility is assumed with respect to all $(\lambda,\kappa)$-almost-geodesics, $\lambda\geq 1,\kappa>0$ (see [17] for the definition). The proof of the above theorem follows with minor modifications from that of [6, Theorem 6.5]. Theorem 1.4 implies that the identity map on a complete Kobayashi hyperbolic domain $\Omega\subsetneq\mathcal{X}$ that satisfies geodesic visibility property extends continuously from $\overline{\Omega}^{\mathsf{k}}$ to $\overline{\Omega}^{\rm End}$.
1.1. Dynamics of holomorphic maps
We now turn to the second theme of this paper, which centers on the asymptotic behavior of iterates of holomorphic self-maps. In particular, we explore analogues of the classical Wolff lemma and Denjoy–Wolff theorem in the setting of complete Kobayashi hyperbolic domains.
To set the stage, let $\xi\in\partial\overline{\mathcal{X}}^{\mathsf{k}}:=\overline{\mathcal{X}}^{\mathsf{k}}\setminus i_{\mathsf{k}}(\mathcal{X})$ be a metric boundary point, and fix a point $p\in\mathcal{X}$. Denote by $h_{p,\,\xi}$ the unique horofunction representing $\xi$ that vanishes at $p$ (see Definition 2.4). The horoball with pole $p$, center $\xi$, and radius $R>0$ is defined as
$H_{p}(\xi,\,R):=\left\{x\in\mathcal{X}\,:\,h_{p,\,\xi}(x)<\tfrac{1}{2}\log R\right\}.$
When $\mathcal{X}$ is a bounded convex domain in $\mathbb{C}^{n}$, the horofunctions $h_{p,\,\xi}$ are quasi-convex, and therefore, $H_{p}(\xi,\,R)$ is a convex open subset of $\mathcal{X}$.
Now let $\Omega\subsetneq\mathcal{X}$ be a Kobayashi hyperbolic domain in a complex manifold $\mathcal{X}$. We consider its end-point compactification $\overline{\Omega}^{\mathrm{End}}$, (see Subsection 2.1), which we assume to be sequentially compact. Given a point $p\in\Omega$, following Marco Abate, we define the big and small horospheres centered at $x\in\partial\overline{\Omega}^{\mathrm{End}}:=\overline{\Omega}^{\mathrm{End}}\setminus\Omega$, with pole $p$ and radius $R>0$, respectively, by:
$H^{b}_{p}(x,R)$ :=$:=\Big{\{}z\in\Omega:\liminf_{w\to x}\left(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\right)<\tfrac{1}{2}\log R\Big{\}},$
$H^{s}_{p}(x,R)$ :=$:=\Big{\{}z\in\Omega:\limsup_{w\to x}\left(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\right)<\tfrac{1}{2}\log R\Big{\}}.$
To relate these horospheres to horoballs, we define, for any $x\in\partial\overline{\Omega}^{\mathrm{End}}$,
$\mathscr{H}(x):=\left\{\xi\in\overline{\Omega}^{\mathsf{k}}\setminus i_{\mathsf{k}}(\Omega)\,:\,\exists(x_{n})\subset\Omega\text{ such that }x_{n}\to x\text{ and }i_{\mathsf{k}}(x_{n})\to\xi\right\}.$
The significance of $\mathscr{H}(x)$ is highlighted by Lemma 4.1, that shows that the big horosphere $H^{b}_{p}(x,R)$ is the union of the horoballs $H_{p}(\xi,R)$ for all $\xi\in\mathscr{H}(x)$, while the small horosphere $H^{s}_{p}(x,R)$ is their intersection.
For bounded convex domains, not only is $H_{p}(\xi,R)$ convex, but its closure also contains any $x\in\partial\overline{\Omega}^{\mathrm{End}}$ such that $\xi\in\mathscr{H}(x)$. This fact provides an explanation for a result due to Abate–Raissy [3] that $\overline{H^{b}_{p}(x,R)}$ is star-shaped with respect to $x$. Building on this framework, and drawing motivation from works of Abate, Karlsson [16], we present a horoball version of Wolff’s lemma adapted to this context.
Theorem 1.5.
Let $\Omega\subsetneq\mathcal{X}$ be a complete Kobayashi hyperbolic domain of a complex manifold $\mathcal{X}$, and suppose that
$\lim_{(z,w)\to(x,y)}\mathsf{k}_{\Omega}(z,w)=\infty\ \ \ \text{for any $x\neq y\in\partial\Omega$}.$ (1.1)
Let $f:\Omega\longrightarrow\Omega$ be a holomorphic map such that $(f^{n})_{n\geq 1}$ is compactly divergent. Then there exist $x\in\partial\overline{\Omega}^{\rm End}$, and $\xi\in\mathscr{H}(x)$ such that, for every $n\in\mathbb{N}$,
$f^{n}(H_{p}(\xi,R))\subset H^{b}_{p}(x,R)\ \ \ \text{for all $R>0$}.$ (1.2)
In particular, for every $n\in\mathbb{N}$, we also have
$f^{n}(H^{s}_{p}(x,R))\subset H^{b}_{p}(x,R)\ \ \ \text{for all $R>0$}.$ (1.3)
The main feature of the above theorem is the inclusion (1.2). In general, small horospheres could be empty for sufficiently small values of $R$, even for very nice domains, see e.g., the case of a slit disk in [14, Example 5.2]. However, horoballs are never empty. This distinction allows for a sharper localization of the target set of a holomorphic self-map, as shown in Lemma 4.7. This, in turn, plays a crucial role in establishing a Denjoy–Wolff theorem.
For compactly approximable domains—such as bounded convex domains—Abate established a Wolff-type lemma (i.e., (1.3)) for fixed-point free holomorphic self-maps by exploiting the interior geometry of the domain. Additionally, Abate [2, Theorem 3.5] proved (1.3) for general domains $\Omega\Subset\mathbb{C}^{n}$ under the assumption that $\Omega$ has a simple boundary, instead of requiring condition (1.1). It is known that the simple boundary condition implies (1.1); see, for instance, [9, Appendix A]. Domains that exhibit visibility with respect to Kobayashi geodesics also satisfy (1.1). However, in general, (1.1) appears to be weaker than both the simple boundary condition and geodesic visibility. We now present the Denjoy–Wolff theorem alluded to above.
Theorem 1.6.
Let $\Omega\subsetneq\mathcal{X}$ be a complete Kobayashi hyperbolic domain that satisfies condition (1.1). Moreover, suppose that for a fixed point $p\in\Omega$ and every $x\in\partial\overline{\Omega}^{\mathrm{End}}$,
$\bigcap_{R>0}\overline{H^{b}_{p}(x,R)}$
is a singleton set. Then, for any holomorphic self-map $f:\Omega\longrightarrow\Omega$, one of the following holds:
  • $(a)$
    For each $z\in\Omega$, the orbit $\{f^{n}(z):n\in\mathbb{N}\}$ is relatively compact in $\Omega$;
  • $(b)$
    There exists a point $y\in\partial\overline{\Omega}^{\mathrm{End}}$ such that $f^{n}\to y$ in the compact-open topology.
By [8, Proposition 2.4], complete hyperbolic domains with the geodesic visibility property satisfy condition (1.1). Moreover, they also satisfy
$\overline{H^{b}_{p}(x,R)}\cap\partial\overline{\Omega}^{\mathrm{End}}=\{x\}\quad\text{for all }R>0,$
as shown in [14, Proposition 1.2]. (This proposition is about bounded domains but the same proof can be adapted for end compactification.) Thus, Theorem 1.6 applies to such domains. For strictly convex domains $\Omega\Subset\mathbb{C}^{n}$, both conditions are verified in [3, Lemmas 3, 5], so the theorem holds in this case as well. Finally, for bounded convex domains, a result of Abate [1, Theorem 2.4.20] implies that part $(a)$ could be replaced by the fact that $f$ has a fixed point.
The Denjoy–Wolff theorem for strictly convex domains was first proved by Budzyńska [12]. Later, Abate–Raissy [3] gave a simpler proof. Their proof relies heavily on convexity, e.g., small horospheres are always nonempty, the target set lies in a linear part of the boundary, and the intersection of all big horospheres centered at a point is a singleton. Motivated by this, we asked whether the last of these — being a metric condition — alone suffices for a Denjoy–Wolff-type result under the condition (1.1), without any geometric assumptions. Our theorem confirms this, potentially allowing applications to broader settings, such as hyperbolic Riemann surfaces embedded in ambient Riemann surfaces.
In the same paper, for a bounded convex domain $D\subset\mathbb{C}^{d}$ and a holomorphic self-map $f$ of $D$, Abate–Raissy [3, Lemma 8] established the existence of an invariant horoball under $f$. Using this result, one can locate the limit set of $f$ — assuming $f$ is fixed-point free — purely in terms of horoballs. Then, adapting the proof of Theorem 1.6 to this setting yields the following result, which is a stronger version of Theorem 1.6 in the context of convex domains.
Proposition 1.7.
Let $\Omega\Subset\mathbb{C}^{d}$ be a bounded convex domain that satisfies condition (1.1). Given $p\in\Omega$, suppose that for every $\xi\in\partial\overline{\Omega}^{\mathsf{k}}$,
$\bigcap_{R>0}\overline{H_{p}(\xi,R)}$ (1.4)
is a singleton. Then for every holomorphic map $f\colon\Omega\to\Omega$, either $f$ has a fixed point in $\Omega$, or there exists a point $y\in\partial\Omega$ such that $f^{n}\to y$ uniformly on compact subsets of $\Omega$.
In a recent preprint, Bracci–Ökten [10, Proposition 4.7] showed that any bounded convex domain with simple boundary satisfies the conclusion of Proposition 1.7 if all sequential horospheres intersect the boundary at exactly one point. This also follows from Proposition 1.7, once we observe that their condition implies
$\overline{H_{p}(\xi,R)}\cap\partial\Omega$
is a singleton for every $R>0$. However, in general, Proposition 1.7 is stronger, as it yields the same conclusion under a weaker hypothesis. The proofs of Theorem 1.5, Theorem 1.6, and Proposition 1.7 are presented in Section 4.
When $(\Omega,\mathsf{k}_{\Omega})$ is Gromov hyperbolic, the identity map extends continuously from $\overline{\Omega}^{\mathsf{k}}$ to the Gromov compactification $\overline{\Omega}^{G}$; see [20]. Since condition (1.1) is built into the structure of Gromov boundaries, adapting the method of Theorem 1.5 could yield a Wolff-type lemma in this setting and may lead to a Denjoy–Wolff theorem for Gromov hyperbolic spaces. This approach suggests a sufficient condition for such a theorem in terms of Busemann horospheres associated with geodesic rays, partially generalizing a recent result of Bracci–Benini [5, Theorem 1.4]. We hope to develop this direction in our forthcoming work.
We now move to the third theme of this paper: namely the continuous extension of the identity map from $\overline{\Omega}^{\mathsf{k}}$ to $\overline{\Omega}^{\rm End}$ and vice-versa. In this direction, we have the following result.
Theorem 1.8.
Let $\Omega\subsetneq\mathcal{X}$ be a Kobayashi hyperbolic domain in a complex manifold $\mathcal{X}$. Then there exists a continuous surjective map $F:\overline{\Omega}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\rm End}$ such that
$F\circ i_{\mathsf{k}}={i}_{\Omega}\ \ \ \text{on $\Omega$}$ (1.5)
if and only if $\mathscr{H}(x)\cap\mathscr{H}(y)=\emptyset$ for all $x\neq y\in\partial\overline{\Omega}^{\rm End}$.
The proof of the above theorem is given in Section 5.1 in which we also provide a sufficient condition (Proposition 5.2) for the existence of $F$, namely:
$\bigcap_{R>0}\overline{H^{b}_{p}(x,R)}=\{x\}$
for each $x\in\partial\overline{\Omega}^{\mathrm{End}}$, which holds, for instance, in domains with geodesic visibility property or bounded strictly convex domains. A necessary condition, for the existence of $F$ as above, is that all geodesic rays land at the boundary, see Proposition 5.1. In the case of Gromov hyperbolicity, Bharali–Zimmer [7, Theorem 1.11] showed that extendability of the identity map from $\overline{\Omega}^{G}$ to $\overline{\Omega}^{\mathrm{End}}$ implies geodesic visibility, and the converse follows from ideas in [8, Theorem 3.3]. Thus, visibility characterizes extendability in the Gromov setting, whereas for the metric compactification, the corresponding condition appears weaker than visibility (cf. Theorem 1.4).
For the converse, the map $i_{\mathsf{k}}:\overline{\Omega}^{\mathrm{End}}\longrightarrow\overline{\Omega}^{\mathsf{k}}$ extends continuously if and only if $\mathscr{H}(x)$ is a singleton for each $x\in\partial\overline{\Omega}^{\mathrm{End}}$, which is equivalent to the existence of the limit
$\lim_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big{)}$
for every $z\in\Omega$. The equivalence between the continuous extension and the existence of the above limit is well known in the context of bounded domains in $\mathbb{C}^{n}$. For completeness, we record this fact in its most general form as Theorem 5.3.
Concluding Remarks. Recall that all strictly convex domains satisfy the Denjoy–Wolff theorem. Moreover, as observed above, in any strictly convex domain, every geodesic ray lands at a boundary point. Motivated by a recent work of Bracci–Benini [5], we pose the following question:
Question 1.
Does there exist $\Omega\Subset\mathbb{C}^{d}$, a bounded strictly convex domain, that does not satisfy the geodesic visibility property?
This question is also related to Question 3.10 in the recent preprint by Bracci–Ökten [10], where the authors ask whether there exist bounded convex domains that are not visibility domains, but for which all geodesic rays land.
We also observe that all strictly convex domains satisfy a logarithmic estimate for the Kobayashi distance, as established in [8, Proposition 4.1]. Therefore, any potential counterexample to the visibility property within the class of strictly convex domains must have a boundary that is not Dini-smooth. In particular, such a domain would need to exhibit relatively low boundary regularity.
2. Preliminaries
In this section, we collect certain preliminary notions, definitions, and auxiliary lemmas that will be useful in the upcoming sections.
2.1. End Compactification
Throughout this paper, we consider a subdomain $\Omega\subsetneq\mathcal{X}$ in a complex manifold $\mathcal{X}$. The end compactification $\overline{\Omega}^{\rm End}$ of $\overline{\Omega}$ in $\mathcal{X}$ is defined as the set
$\overline{\Omega}^{\rm End}:=\overline{\Omega}\,\sqcup\,\mathcal{E},$
where $\mathcal{E}$ is the set of all ends of $\overline{\Omega}$. An end is defined as follows: fix a compact exhaustion $(K_{j})_{j\geq 1}$ of $\overline{\Omega}$ such that $K_{j}\subset{\rm int}(K_{j+1})$ for each $j$. An end $e$ is a sequence $(F_{j})_{j\geq 1}$ where each $F_{j}$ is a connected component of $\overline{\Omega}\setminus K_{j}$, and $F_{j+1}\subset F_{j}$ for all $j\geq 1$. Note that if $\Omega\Subset\mathcal{X}$, then $\mathcal{E}=\emptyset$. We now describe a topology on $\overline{\Omega}^{\rm End}$. For $x\in\overline{\Omega}$, a neighborhood basis is given by all neighborhoods of $x$ in $\mathcal{X}$ intersected with $\overline{\Omega}$. For an end $e=(F_{j})_{j\geq 1}$, the neighborhood basis is given by the family $(\widehat{F}_{j})_{j\geq 1}$ defined as
$\widehat{F}_{j}:=F_{j}\,\sqcup\,\left\{f\in\mathcal{E}\,\big{|}\,f=(G_{\nu})_{\nu\geq 1}\text{ with }G_{\nu}=F_{\nu}\ \text{for }\nu=1,\ldots,j\right\}.$
With this topology, $\overline{\Omega}^{\rm End}$ becomes a first-countable Hausdorff topological space, and the resulting topology does not depend on the choice of compact exhaustion. In general, $\overline{\Omega}^{\rm End}$ need not be compact or even sequentially compact. However, when $\overline{\Omega}$ is locally connected, compactness of $\overline{\Omega}^{\rm End}$ follows. A sufficient condition for sequential compactness is provided in [13, Result 2.1]; in particular, any form of visibility property ensures sequential compactness. Throughout this paper, we shall assume that $\overline{\Omega}^{\rm End}$ is sequentially compact. We now present a lemma that is needed later. In what follows, $h$ will denote a hermitian metric on $\mathcal{X}$, and $d_{h}$ will be the distance induced by $h$.
Lemma 2.1.
Let $\Omega\subsetneq\mathcal{X}$ be a domain in a complex manifold $\mathcal{X}$ such that $\overline{\Omega}^{\mathrm{End}}$ is sequentially compact. Let $(x_{n})_{n\geq 1}\subset\overline{\Omega}$ be a sequence that is not relatively compact in $\mathcal{X}$. Suppose there exists a compact set $K\subset\overline{\Omega}$ and a connected component $F_{0}$ of $\overline{\Omega}\setminus K$ such that $x_{n}\in F_{0}$ for all $n$. Then $x_{n}\in\mathrm{int}(F_{0})$ for all but finitely many $n$.
Proof.
Let $\{F_{0}\}\cup\{F_{\lambda}:\lambda\in\Lambda\}$ be the collection of all connected components of $\overline{\Omega}\setminus K$, where each $F_{\lambda}$ is closed in $\overline{\Omega}\setminus K$.
Claim. If $x_{n}\notin\mathrm{int}(F_{0})$ for some $n$, then for any $r>0$, the ball $B_{d_{h}}(x_{n},r)$ intersects infinitely many components $F_{\lambda}\neq F_{0}$.
Suppose not. Then there exists $r>0$ such that $B_{d_{h}}(x_{n},r)$ intersects only finitely many components other than $F_{0}$, say $F_{\lambda_{1}},\ldots,F_{\lambda_{k}}$. Now, for every $\epsilon\in(0,r)$, the ball $B_{d_{h}}(x_{n},\epsilon)$ intersects at least one of these $F_{\lambda_{i}}$’s, implying that $x_{n}$ is a limit point of $F_{\lambda_{1}}\cup\cdots\cup F_{\lambda_{k}}$, which is a finite union of closed sets. Hence, $x_{n}$ must belong to one of them, contradicting the assumption that $x_{n}\in F_{0}$. This proves the claim.
Now suppose that $x_{n}\notin\mathrm{int}(F_{0})$ for infinitely many $n$. Without loss of generality, assume $x_{n}\notin\mathrm{int}(F_{0})$ for all $n$. Fix $r>0$. For each $n$, use the claim to choose a point $w_{n}\in B_{d_{h}}(x_{n},r)\cap F_{\lambda_{n}}$ such that $\lambda_{n}\in\Lambda$ and $F_{\lambda_{n}}\notin\{F_{\lambda_{1}},\ldots,F_{\lambda_{n-1}}\}$. This is possible since each $B_{d_{h}}(x_{n},r)$ intersects infinitely many components. Then $(w_{n})$ is not relatively compact in $\mathcal{X}$. We now show that $(w_{n})$ has no convergent subsequence in $\overline{\Omega}^{\mathrm{End}}$. Let $(K_{j})_{j\geq 1}$ be a compact exhaustion of $\overline{\Omega}$ with $K_{1}=K$. Since each $w_{n}$ lies in a different component of $\overline{\Omega}\setminus K_{1}$, no end can capture infinitely many $w_{n}$, and there is no subsequence converging to a point in $\overline{\Omega}^{\mathrm{End}}$. This contradicts the sequential compactness of $\overline{\Omega}^{\mathrm{End}}$, completing the proof. ∎
2.2. Metric compactification
In this subsection, we review the construction of the metric compactification for a general metric space. To facilitate its use later, we also include a few foundational results — both statements and proofs — that are standard but not always readily referenced.
Given a metric space $(X,d)$, let $C(X)$ be the space of all real-valued continuous functions on $X$. We equip $C(X)$ with the topology of local uniform convergence. Let $C_{*}(X)$ be the quotient space of $C(X)$ by the subspace of constant functions. Consider the canonical projection $\pi:C(X)\longrightarrow C_{*}(X)$ defined by
$\pi(f):=[\,f\,]:=\left\{f+c:c\in\mathbb{R}\right\}.$
[…]
Lemma 2.2.
Let $p\in X$ be given. Then the map $T_{p}:C_{*}(X)\longrightarrow C_{p}(X)$ given by $T_{p}([f]):=f-f(p)$ is a homeomorphism.
The proof is elementary and is therefore omitted.
Consider the map $i_{d}:X\longrightarrow C_{*}(X)$ defined by $i_{d}(x):=[\,d_{x}(\boldsymbol{\cdot})\,]$, where $d_{x}:X\to\mathbb{R}$ is the continuous function defined by $d_{x}(z):=d(x,z)$. When we fix a point $p\in X$, $T_{p}\circ i_{d}(x)$ is the continuous function $d(x,\cdot)-d(x,p)$ for each $x\in X$.
Lemma 2.3.
Let $(X,d)$ be a metric space. Then the map $i_{d}:X\longrightarrow C_{*}(X)$ is an injective continuous map.
Proof.
It is sufficient to show that for a fixed $p\in X$, $T_{p}\circ i_{d}:X\longrightarrow C_{p}(X)$ is one-one and continuous. Suppose $T_{p}\circ i_{d}(x)=T_{p}\circ i_{d}(y)$ for some $x,y\in X$. Therefore $d_{x}(\boldsymbol{\cdot})-d_{x}(p)=d_{y}(\boldsymbol{\cdot})-d_{y}(p)$ on $X$. By putting $x$ and $y$ separately in the last equation, we get $d(x,y)=0$, i.e. $x=y$. Hence $T_{p}\circ i_{d}$ is one-one. Now we show that $T_{p}\circ i_{d}$ is continuous. Let $(x_{n})$ be a sequence convergent to $x$ in $X$. Then
$|T_{p}\circ i_{d}(x_{n})-T_{p}\circ i_{d}(x)|=|d_{x_{n}}(\boldsymbol{\cdot})-d(p,x_{n})-d_{x}(\boldsymbol{\cdot})+d(p,x)|\leq 2\,d(x_{n},x).$
Therefore, $T_{p}\circ i_{d}(x_{n})$ converges to $T_{p}\circ i_{d}(x)$ uniformly in $C_{p}(X)$. Hence $T_{p}\circ i_{d}$ is continuous, from which the result follows. ∎
Definition 2.4.
Let $(X,d)$ be a metric space. The metric compactification denoted by $\overline{X}^{d}$ of $X$ is the closure of $i_{d}(X)$ in $C_{*}(X)$ and the metric boundary of $X$ is the set
$\partial\overline{X}^{d}:=\overline{X}^{d}\setminus i_{d}(X).$
Given $\xi\in\partial\overline{X}^{d}$, there exists a function $h_{\xi}\in C(X)$, unique up to an additive constant, such that $[\,h_{\xi}\,]=\xi$. The function $h_{\xi}$ is called a horofunction corresponding to $\xi$. In particular, if we fix a point $p\in X$, then there is a unique such function $h_{p,\,\xi}$ which vanishes at the point $p$. Given $R>0$, the set
$H_{p}(\xi,R):=\left\{x\in X\,:\,h_{p,\,\xi}(x)<\tfrac{1}{2}\log R\right\}$
is called the horoball of radius $R$ centered at $\xi$ with pole $p$.
Lemma 2.5.
Let $(X,d)$ be a metric space. Then the metric compactification $\overline{X}^{d}$ of $X$ is compact.
Proof.
Fix a point $p\in X$. Using Lemma 2.2, we can identify $\overline{X}^{d}$ as the closure of the family of functions $\{T_{p}\circ i_{d}(x):x\in X\}$. Note that
$|T_{p}\circ i_{d}(x)(y_{1})-T_{p}\circ i_{d}(x)(y_{2})|\leq d(y_{1},y_{2})\ \ \ \forall x,y_{1},y_{2}\in X$
Therefore, the family $\{T_{p}\circ i_{d}(x):x\in X\}$ is equicontinuous. Also note that for any $y\in X$
$|T_{p}\circ i_{d}(x)(y)|\leq d(p,y).$
So the family $\{T_{p}\circ i_{d}(x):x\in X\}$ is uniformly bounded. Since $X$ is locally compact, by Ascoli’s Theorem [18, Theorem 47.1], the family $\{T_{p}\circ i_{d}(x):x\in X\}$ is relatively compact in $C_{p}(X)$. Hence $\overline{X}^{d}$ is compact. ∎
Finally, given $\epsilon>0$, we recall that a map $\sigma:I\longrightarrow(X,d)$ is called a $(1,\epsilon)$-quasi-geodesic if
$|t_{1}-t_{2}|-\epsilon\leq d(\sigma(t_{1}),\sigma(t_{2}))\leq|t_{1}-t_{2}|+\epsilon\quad\forall t_{1},t_{2}\in I,$
where $I\subset\mathbb{R}$ is an interval. The following lemma follows easily from the definition of $(1,\epsilon)$-quasi-geodesic.
Lemma 2.6.
Let $(X,d)$ be a metric space and let $x,y,z\in X$ be points that lie on $(1,\epsilon)$-quasi-geodesic. Then
$d(x,z)\leq d(x,y)+d(y,z)\leq d(x,z)+3\epsilon$
3. Metric compactification of a hyperbolic complex manifold
In this section, we present the proofs of Theorem 1.2, Proposition 1.3 and Theorem 1.4. For a complex manifold $\mathcal{X}$, given $p,z\in\mathcal{X}$, in what follows, we shall use the notation $\phi_{z,\,p}(\boldsymbol{\cdot})$ to denote the function $T_{p}\circ i_{\mathsf{k}}(z)=\mathsf{k}_{\mathcal{X}}(\boldsymbol{\cdot},z)-\mathsf{k}_{\mathcal{X}}(p,z)$. We shall also denote $\phi_{z,\,p}$ simply by $\phi_{z}$ if $p$ is fixed throughout the discussion. The following result was already mentioned in the introduction, and is needed in our proof of Theorem 1.2.
Result 3.1 (paraphrasing [17, Proposition 2.8]).
Let $\mathcal{X}$ be a Kobayashi hyperbolic complex manifold equipped with the Kobayashi distance $\mathsf{k}_{\mathcal{X}}$. Then, given $\epsilon>0$, for any two points $z,w\in\mathcal{X}$, there exists a $(1,\epsilon)$-quasi-geodesic connecting them.
We are now ready to present
The proof of Theorem 1.2.
Fix a point $p\in\mathcal{X}$. Appealing to Lemma 2.5 and Lemma 2.3, it is sufficient to show that $T_{p}\circ i_{\mathsf{k}}$ is an open map onto its image in $C_{p}(\mathcal{X})$. Let $U$ be an open set in $\mathcal{X}$. Observe that
$T_{p}\circ i_{\mathsf{k}}(U)=\Big{\{}\mathsf{k}_{\mathcal{X}}(\boldsymbol{\cdot},z)-\mathsf{k}_{\mathcal{X}}(p,z):z\in U\Big{\}}.$
We shall show the complement of $T_{p}\circ i_{\mathsf{k}}(U)$ in $T_{p}\circ i_{\mathsf{k}}(\mathcal{X})$ is closed. Consider the sequence given by $\phi_{z_{n}}(\boldsymbol{\cdot}):=\mathsf{k}_{\mathcal{X}}(\boldsymbol{\cdot},z_{n})-\mathsf{k}_{\mathcal{X}}(p,z_{n})$, where $(z_{n})\subset\mathcal{X}\setminus U$, that converges locally uniformly to $\phi_{w}(\boldsymbol{\cdot})$ for some $w\in\mathcal{X}$. We now show that $w\in\mathcal{X}\setminus U$. Two cases arise:
Case 1: The sequence $(z_{n})$ has a subsequence that converges to $z_{0}\in\mathcal{X}$. Assume, by reindexing if necessary, that $z_{n}\to z_{0}$. Then, as observed in Lemma 2.3, $\phi_{z_{n}}(\boldsymbol{\cdot})$ converges locally uniformly to $\phi_{z_{0}}(\boldsymbol{\cdot})$. By injectivity of $T_{p}\circ i_{\mathsf{k}}$, we get $w=z_{0}$. Since $\mathcal{X}\setminus U$ is closed, $w=z_{0}\in\mathcal{X}\setminus U$ and we are done.
Case 2: The sequence $(z_{n})$ has no limit point in $\mathcal{X}$. Since $\big{(}\mathcal{X},\mathsf{k}_{\mathcal{X}}\big{)}$ is locally compact, there exists an $r>0$ such that $\overline{B_{\mathsf{k}}(w,r)}$ is compact in $\big{(}\mathcal{X},\mathsf{k}_{\mathcal{X}}\big{)}$. Here, ${B_{\mathsf{k}}(w,r)}=\{z\in\mathcal{X}:\mathsf{k}_{\mathcal{X}}(z,w)<r\}$. The sequence $(z_{n})$ has only finitely many points in the closed ball $\overline{B_{\mathsf{k}}(w,r)}$. Therefore, we may assume that for each $n\geq 1$, $\mathsf{k}_{\mathcal{X}}(z_{n},w)>r$. Let $\epsilon=r/3$, and choose $\sigma_{n}$, a $(1,\epsilon)$-quasi-geodesic connecting $z_{n}$ and $w$, by applying Result 3.1. Let $w_{n}\in\partial B_{\mathsf{k}}(w,r)$ be such that $w_{n}\in{\rm Image}(\sigma_{n})$. Since $\partial B_{\mathsf{k}}(w,r)$ is compact, by passing to a subsequence and reindexing, we can assume that $w_{n}\to w_{r}$ for some $w_{r}\in\partial B_{\mathsf{k}}(w,r)$. Now, since $\phi_{z_{n}}\to\phi_{w}$ locally uniformly, we have
$\lim_{n\to\infty}\phi_{z_{n}}(w_{n})=\phi_{w}(w_{r})=\mathsf{k}_{\mathcal{X}}(w,w_{r})-\mathsf{k}_{\mathcal{X}}(w,p)=r-\mathsf{k}_{\mathcal{X}}(w,p).$ (3.1)
On the other hand, for each $n\in\mathbb{N}$, we have $\phi_{z_{n}}(w_{n})=\mathsf{k}_{\mathcal{X}}(z_{n},w_{n})-\mathsf{k}_{\mathcal{X}}(z_{n},p)$. Since the points $w$, $w_{n}$ and $z_{n}$ lie on a $(1,\epsilon)$-quasi-geodesic, by Lemma 2.6, for each $n\in\mathbb{N}$, we have
$\mathsf{k}_{\mathcal{X}}(w,w_{n})+\mathsf{k}_{\mathcal{X}}(w_{n},z_{n})\leq\mathsf{k}_{\mathcal{X}}(w,z_{n})+3\epsilon\implies\mathsf{k}_{\mathcal{X}}(w_{n},z_{n})\leq\mathsf{k}_{\mathcal{X}}(w,z_{n})-r+3\epsilon\ \ \ \forall n\in\mathbb{N}.$
Therefore, for every $n\in\mathbb{N}$,
$\mathsf{k}_{\mathcal{X}}(w_{n},z_{n})-\mathsf{k}_{\mathcal{X}}(z_{n},p)\leq\mathsf{k}_{\mathcal{X}}(w,z_{n})-\mathsf{k}_{\mathcal{X}}(z_{n},p)-r+3\epsilon\implies\phi_{z_{n}}(w_{n})\leq\phi_{z_{n}}(w)-r+3\epsilon$
for all $n$. Taking $n\to\infty$, and using (3.1), we get
$r-\mathsf{k}_{\mathcal{X}}(w,p)\leq-\mathsf{k}_{\mathcal{X}}(w,p)-r+3\epsilon$
which implies that $2r\leq 3\epsilon=r$, which is absurd. Hence this case does not arise. ∎
Remark 3.2.
Note that it follows from the proof of Theorem 1.2 that $\xi\in\partial\overline{\mathcal{X}}^{\mathsf{k}}:=\overline{\mathcal{X}}^{\mathsf{k}}\setminus i_{\mathsf{k}}(\mathcal{X})$, i.e., $\xi$ is a metric boundary point if and only if there exists a sequence $(z_{\nu})_{\nu\geq 1}\subset\mathcal{X}$ such that $i_{\mathsf{k}}(z_{\nu})\to\xi$ and $(z_{\nu})_{\nu\geq 1}$ does not have a limit point in $\mathcal{X}$.
The proof of the above theorem could be adapted to any locally compact metric space in which every pair of points can be joined by a $(1,\epsilon)$-quasi-geodesic for any $\epsilon>0$.
Focusing our attention on a proper geodesic metric space $(X,d)$, where $(\overline{X}^{d},i_{d})$ forms a compactification, we now show that it is, in fact, a good compactification in the sense of Bharali–Zimmer [6], defined as follows.
Definition 3.3.
[…]
$\lim_{n\to\infty}\iota(\sigma_{n}(a_{n}))=\lim_{n\to\infty}\iota(\sigma_{n}(b_{n}))\in{X}^{*}\setminus\iota(X)$
we have
$\liminf_{n\to\infty}d(o,\sigma_{n})=\infty$
for any $o\in X$.
We require a few foundational lemmas due to Webster–Winchester [20] to establish our main result. Since their proofs are elementary and the results fundamental, we include self-contained proofs adopted to our setting. The first lemma [20, Lemma 4.1] essentially says that horoballs of arbitrarily small radius are nonempty. In our work [14, Lemma 3.2], we independently observed that the big horospheres are always nonempty. The proof of the lemma below follows the idea presented in the latter work.
Lemma 3.4.
Let $(X,d)$ be a proper, geodesic metric space and let $h\in C(X)$ be a horofunction. Then for any $R<0$, there exists $x\in X$ such that $h(x)<R$.
Proof.
Fix a point $p\in X$, then there exists a sequence $(x_{n})_{n\geq 1}\subset X$ such that $(d(\cdot,x_{n})-d(p,x_{n}))_{n\geq 1}$ converges locally uniformly to $h(\cdot)-h(p)$. Note, as noted in Remark 3.2, $(x_{n})_{n\geq 1}$ cannot have a limit point in $X$. Since $X$ is a proper metric space, this implies $\lim_{n\to\infty}d(p,x_{n})=\infty$. For $s>0$, consider the ball $B_{d}(p,s)$. Observe, we may assume that $d(p,x_{n})>s$ for all $n\geq 1$. For each $n$, choose $b_{n}\in\partial B_{d}(p,s)$ such that
$d(p,x_{n})=d(p,b_{n})+d(b_{n},x_{n})\quad\forall n\in\mathbb{N}.$ (3.2)
By passing to a subsequence, we may assume $b_{n}\to b_{s}$ as $n\to\infty$, for some $b_{s}\in\partial B_{d}(p,s)$. Therefore, given $\epsilon>0$, there exists $n_{0}\in\mathbb{N}$ such that $d(b_{s},b_{n})<\epsilon$ for all $n\geq n_{0}$. Therefore,
$d(b_{s},x_{n})\leq d(b_{s},b_{n})+d(b_{n},x_{n})<\epsilon+d(b_{n},x_{n})\quad\forall n\geq n_{0}.$ (3.3)
Using (3.2) and (3.3),we get
$d(b_{s},x_{n})-d(p,x_{n})<-s+\epsilon\quad\forall n\geq n_{0}.$
Taking $n\to\infty$, we get $h(b_{s})-h(p)\leq-s+\epsilon$ for all $s>0$, $\epsilon>0$. Therefore, $h(b_{s})<-s/2+h(p)$ for all $s>0$. Now by choosing $s>2(h(p)-2R)$, we get the result. ∎
Given points $x,y\in(X,d)$, the Gromov product of $x,y$ with respect to a point $p\in X$, denoted by $\langle x,y\rangle_{p}$, is defined by
$\langle x,y\rangle_{p}:=\frac{d(x,p)+d(y,p)-d(x,y)}{2}.$
We now present the next lemma.
Lemma 3.5.
Let $(X,d)$ be a metric space. Fix a point $p\in X$, then for any $x,y,z\in X$, we have
$\langle x,y\rangle_{p}\geq-\frac{\big{(}T_{p}\circ i_{d}(x)(z)+T_{p}\circ i_{d}(y)(z)\big{)}}{2},$
where $T_{p}$ is as in Lemma 2.2.
Proof.
Let $z$ be any point in $X$ then
$2\langle x,y\rangle_{p}$ $=d(x,p)+d(y,p)-d(x,y)\geq d(x,p)+d(y,p)-d(x,z)-d(z,y)$
$=-\big{(}d(x,z)-d(x,p)\big{)}-\big{(}d(z,y)-d(y,p)\big{)}$
$=-T_{p}\circ i_{d}(x)(z)-T_{p}\circ i_{d}(y)(z),$
from which the result follows. ∎
Lemma 3.6.
Let $(X,d)$ be a proper, geodesic metric space. Let $(x_{n})_{n\geq 1}\subset X$ be a sequence such that $(i_{d}(x_{n}))_{n\geq 1}\subset C_{*}(X)$ converges to a horofunction. Then, for any $p\in X$, we have
$\lim_{n\to\infty}\big{<}x_{n},x_{m}\big{>}_{p}=\infty.$
Proof.
Fix $p\in X$ then $T_{p}\circ i_{d}(x_{n})\longrightarrow h\in C(X)$ locally uniformly and $h$ is a horofunction. Let $M>0$ be given then by Lemma 3.4, we can find $z\in X$ such that $h(z)<-M$. Since $T_{p}\circ i_{d}(x_{n})\longrightarrow h\in C(X,\mathbb{R})$ locally uniformly, there exists $N_{0}\in\mathbb{N}$ such that $T_{p}\circ i_{d}(x_{n})(z)<{-M}/{2}$ for all $n\geq N_{0}$. Therefore, for $n,m\geq N_{0}$, we get
$\big{<}x_{n},x_{m}\big{>}_{p}\geq-\frac{\big{(}T_{p}\circ i_{d}(x_{n})(z)+T_{p}\circ i_{d}(x_{m})(z)\big{)}}{2}\geq\frac{M}{2}$
Hence the result. ∎
The next lemma is in similar vein and easily follows from the above. So we only state the lemma.
Lemma 3.7.
Let $(X,d)$ be a proper, geodesic metric space. Let $(x_{n}),(y_{n})\subset X$ be two sequences such that $(i_{d}(x_{n}))$ and $(i_{d}(y_{n}))$ converge to the same horofunction then
$\lim_{n,m\to\infty}\big{<}x_{n},y_{m}\big{>}_{p}=\infty.$
We now have all the tools to present
The proof of Proposition 1.3.
[…]
Claim. $d\big{(}p,\sigma_{n}([a_{n},b_{n}])\big{)}\geq\langle x_{n},y_{n}\rangle_{p}$.
To see the claim, let $z_{n}\in\sigma_{n}([a_{n},b_{n}])$ be such that $d(p,z_{n})=d\big{(}p,\sigma_{n}([a_{n},b_{n}])\big{)}$. Then, using the triangle inequality and that $\sigma_{n}$ is a geodesic, we have
$2d\big{(}p,\sigma_{n}([a_{n},b_{n}])\big{)}$ $=2d(p,z_{n})$
$\geq(d(p,x_{n})-d(x_{n},z_{n}))+(d(p,y_{n})-d(y_{n},z_{n}))$
$=d(p,x_{n})+d(p,y_{n})-d(x_{n},y_{n}),$
from which the claim follows.
By Lemma 3.7, we have $\lim_{n\to\infty}\big{<}x_{n},y_{n}\big{>}_{p}=\infty$, and therefore by the claim above,
$\liminf_{n\to\infty}d(p,\sigma_{n}[a_{n},b_{n}])=\infty.$
This establishes the result that $(\overline{X}^{d},i_{d})$ is a good compactification of $X$. ∎
We now have all the ingredients to present
The proof of Theorem 1.4.
Note by Proposition 1.3, $(\overline{\mathcal{X}}^{\mathsf{k}},i_{\mathsf{k}})$ is a good compactification of $\mathcal{X}$. Moreover, it follows from Theorem [17, Theorem 1.16] that $\Omega$ is a visibility domain subordinate to $\overline{\Omega}^{\rm End}$. Using exactly the same arguments as in the proof of [6, Theorem 6.5], it follows that for every $\xi\in\partial\overline{\mathcal{X}}^{\mathsf{k}}$, there exists a $x\in\partial\overline{\Omega}^{\rm End}$ such that
$\lim_{i_{\mathsf{k}}(z)\to\xi}F(z)=x.$
Now consider the map $\widetilde{F}:\overline{\mathcal{X}}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\rm End}$ defined by
$\widetilde{F}(\xi):=\begin{cases}F(\xi),&\ \ \ \xi\in\mathcal{X},\\ \lim_{i_{\mathsf{k}}(z)\to\xi}F(z),&\ \ \ \text{if $\xi\in\partial\overline{\mathcal{X}}^{\mathsf{k}}$.}\end{cases}$
It is now routine to check that $\widetilde{F}$ is continuous. ∎
4. Horoballs, a Wolff-type lemma and a Denjoy–Wolff Theorem
In this section, we shall present the proofs of our main results regarding the dynamics of a holomorphic self-map of a Kobayashi hyperbolic domain of a complex manifold. But first we begin with describing the exact relation beween horoballs and the small and big horospheres as mentioned in the introduction.
4.1. Horoballs, small and big horospheres
Let $\Omega\subsetneq\mathcal{X}$ be a Kobayashi hyperbolic domain of a complex manifold $\mathcal{X}$. We shall denote by $(\overline{\Omega}^{\mathsf{k}},i_{\mathsf{k}})$ the metric compactification of the distance space $(\Omega,\mathsf{k}_{\Omega})$. Recall that by Theorem 1.2, $\overline{\Omega}^{\mathsf{k}}$ is a compactification of $\Omega$ as defined in Definition 1.1. We shall also consider $\overline{\Omega}^{\rm End}$, the end-compactification of $\Omega$ as introduced in Subsection 2.1. As remarked in that subsection, we will always assume that $\overline{\Omega}^{\rm End}$ is sequentially compact. Given $x\in\partial\overline{\Omega}^{\rm End}:=\overline{\Omega}^{\rm End}\setminus\Omega$, recall
$\mathscr{H}(x):=\big{\{}\xi\in\overline{\Omega}^{\mathsf{k}}\setminus i_{\mathsf{k}}(\Omega)\,:\,\text{$\exists(x_{n})_{n\geq 1}\subset\Omega$ such that $x_{n}\to x$ and $i_{\mathsf{k}}(x_{n})\to\xi$}\big{\}}.$
Our first lemma precisely describes the relation between the big, small horospheres and the horoballs centered at metric boundary points.
Lemma 4.1.
Given $p\in\Omega$, $R>0$ and $x\in\partial\overline{\Omega}^{\rm End}$, we have
$H^{b}_{p}(x,R)=\bigcup_{\xi\in\mathscr{H}(x)}H_{p}(\xi,R),\quad\text{and}\quad H^{s}_{p}(x,R)=\bigcap_{\xi\in\mathscr{H}(x)}H_{p}(\xi,R),$
where $H^{b}_{p}(x,R),H^{s}_{p}(x,R)$ are the big and small horospheres with center $x$, pole $p$ and radius $R$ respectively.
Proof.
Recall that
$H^{b}_{p}(x,R)=\left\{z\in\Omega\,:\,\liminf_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big{)}<\tfrac{1}{2}\log R\right\}.$
Let $\xi\in\mathscr{H}(x)$ and let $h_{p,\,\xi}$ be the horofunction representing $\xi$ that vanishes at $p$. Then there exists a sequence $(x_{n})\subset\Omega$, $x_{n}\to x$, such that
$h_{p,\,\xi}(\cdot)=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\big{)}.$
Now suppose $z\in H_{p}(\xi,R)$ then
$\liminf_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big{)}\leq\lim_{n\to\infty}(\mathsf{k}_{\Omega}(z,x_{n})-\mathsf{k}_{\Omega}(p,x_{n}))=h_{p,\,\xi}(z)<\tfrac{1}{2}\log R.$
This implies that for each $\xi\in\mathscr{H}(x)$, $H_{p}(\xi,R)\subset H^{b}_{p}(x,R)$. Therefore, we have
$\bigcup_{\xi\in\mathscr{H}(x)}H_{p}(\xi,R)\subset H^{b}_{p}(x,R).$
Now let $z\in H^{b}_{p}(x,R)$, then there exists a sequence $(x_{n})$ such that $x_{n}\to x$ and
$\liminf_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big{)}=\lim_{n\to\infty}(\mathsf{k}_{\Omega}(z,x_{n})-\mathsf{k}_{\Omega}(p,x_{n}))<\tfrac{1}{2}\log R.$
Consider the sequence $(\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(p,x_{n}))$, as $x_{n}\to x$, by the compactness of $\overline{\Omega}^{\mathsf{k}}$, a subsequence of this sequence converges to a horofunction that vanishes at $p$. If we write this function as $h_{p,\,\xi}$ then $[\,h_{p,\,\xi}\,]=\xi\in\mathscr{H}(x)$. The inequality above implies that
$h_{p,\,\xi}(z)<\tfrac{1}{2}\log R$
whence $z\in H_{p}(\xi,R)$. Since $z$ is arbitrary, we are done.
For small horospheres, choose $z\in H^{s}_{p}(x,R)$, where, recall
$H^{s}_{p}(x,R)=\left\{z\in\Omega\,:\,\limsup_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big{)}<\tfrac{1}{2}\log R\right\},$
and let $\xi\in\mathscr{H}(x)$ be given. Consider $h_{p,\,\xi}$ as above and let $(x_{n})\subset\Omega$, $x_{n}\to x$, such that $h_{p,\,\xi}(\cdot)=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\big{)}$. Therefore,
$h_{p,\,\xi}(z)$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(z,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\big{)}\leq\limsup_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big{)}<\tfrac{1}{2}\log R.$
Since $\xi$ is arbitrary, we get that
$H^{s}_{p}(x,R)\subset\bigcap_{\xi\in\mathscr{H}(x)}H_{p}(\xi,R).$
Now let $z\in\Omega$ be such that $z\in H_{p}(\xi,R)$ for every $\xi\in\mathscr{H}(x)$. Now choose a sequence $(x_{n})$ such that $x_{n}\to x$ and
$\limsup_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(p,w)\big{)}=\lim_{n\to\infty}(\mathsf{k}_{\Omega}(z,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})).$
Again, by the compactness of $\overline{\Omega}^{\mathsf{k}}$, a subsequence of $(\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(p,x_{n}))_{n\geq 1}$ converges to a horofunction that vanishes at $p$. If we write this function as $h_{p,\,\xi}$ then $[\,h_{p,\,\xi}\,]=\xi\in\mathscr{H}(x)$. By our assumption $h_{p,\,\xi}(z)<(1/2)\log R$, and the equality above implies that $z\in H^{s}_{p}(x,R)$. ∎
For bounded convex domains, the horofunctions are quasi-convex.
Proposition 4.2.
Let $\Omega\Subset\mathbb{C}^{n}$ be a convex domain. Given $\xi\in\partial\overline{\Omega}^{\mathsf{k}}$, let $x\in\partial\Omega$ be such that $\xi\in\mathscr{H}(x)$. Then, for a fixed $p\in\Omega$, we have
$h_{p,\,\xi}\big{(}tz_{1}+(1-t)z_{2}\big{)}\leq\max\left\{h_{p,\,\xi}(z_{1}),h_{p,\,\xi}(z_{2})\right\},\quad\forall z_{1},z_{2}\in\Omega,\ t\in[0,1].$ (4.1)
In particular, each horoball $H_{p}(\xi,R)$ is convex, and moreover, $x\in\overline{H_{p}(\xi,R)}$.
Proof.
By the definition of $h_{p,\,\xi}$, since $\xi\in\mathscr{H}(x)$, there exists a sequence $(x_{n})$ such that $x_{n}\to x$ and $h_{p,\,\xi}(\cdot)=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\big{)}$. Given $z_{1},z_{2}\in\Omega$, $t\in[0,1]$, and $x_{n}$, the convexity of $\Omega$ implies that $\mathsf{k}_{\Omega}(tz_{1}+(1-t)z_{2},x_{n})\leq\max\left\{\mathsf{k}_{\Omega}(z_{1},x_{n}),\mathsf{k}_{\Omega}(z_{2},x_{n})\right\}$, see e.g., [1, Proposition 2.3.46]. Therefore,
$\mathsf{k}_{\Omega}(tz_{1}+(1-t)z_{2},x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\leq\max\left\{\mathsf{k}_{\Omega}(z_{1},x_{n})-\mathsf{k}_{\Omega}(p,x_{n}),\,\mathsf{k}_{\Omega}(z_{2},x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\right\}.$
Taking limit as $n\to\infty$ in the above equation, (4.1) follows. Clearly this implies that $H_{p}(\xi,R)$ is convex. To show that $x\in\overline{H_{p}(\xi,R)}$, it is enough to show that for every $z\in H_{p}(\xi,R)$, $s\in(0,1)$, $sz+(1-s)x\in H_{p}(\xi,R)$. Write
$\mathsf{k}_{\Omega}(sz+(1-s)x,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})$ $\leq\big{(}\mathsf{k}_{\Omega}(sz+(1-s)x_{n},x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\big{)}$
$+\big{(}\mathsf{k}_{\Omega}(sz+(1-s)x,x_{n})-\mathsf{k}_{\Omega}(sz+(1-s)x_{n},x_{n})\big{)}$
Note that again the convexity of $\Omega$ implies that $\mathsf{k}_{\Omega}(sz+(1-s)x_{n},x_{n})\leq\mathsf{k}_{\Omega}(z,x_{n})$. Also,
$\lim_{n\to\infty}\mathsf{k}_{\Omega}(sz+(1-s)x,x_{n})-\mathsf{k}_{\Omega}(sz+(1-s)x_{n},x_{n})\leq\lim_{n\to\infty}\mathsf{k}_{\Omega}(sz+(1-s)x,sz+(1-s)x_{n})=0.$
Therefore, we have
$\lim_{n\to\infty}\mathsf{k}_{\Omega}(sz+(1-s)x,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})\leq\lim_{n\to\infty}\mathsf{k}_{\Omega}(z,x_{n})-\mathsf{k}_{\Omega}(p,x_{n})<\tfrac{1}{2}\log R.$
Since $s\in(0,1)$ is arbitrary, it follows that $x\in\overline{H_{p}(\xi,R)}$. ∎
The following result — already established by Abate–Raissy [3] — is a direct consequence of Lemma 4.1 and Proposition 4.2.
Corollary 4.3.
Let $\Omega$ be a bounded convex domain. Then for any $p\in\partial\Omega$, $x\in\partial\Omega$ and $R>0$, $H^{s}_{p}(x,R)$ is convex; $\overline{H^{b}_{p}(x,R)}$ is star-shaped with respect to $x$.
4.2. Limit set and target set
In this subsection, we recall the notion of the limit set and target set for a holomorphic self-map. We also prove two auxiliary lemmas that will help us in the proof of our main theorems in the next subsection.
Definition 4.4.
Let $\mathcal{X}$ be a complex manifold and let $\Omega\subsetneq\mathcal{X}$ be a subdomain. Consider the space of continuous functions from $\Omega$ to $\overline{\Omega}^{\rm End}$, together with compact-open topology. Let $f:\Omega\longrightarrow\Omega$ be a holomorphic self-map, or more generally, a self-map that is non-expansive with respect to the Kobayashi distance on $\Omega$. A map $g:\Omega\longrightarrow\overline{\Omega}^{\rm End}$ is called a limit map of the sequence of iterates $(f^{k})$ if there exists a subsequence $\{f^{k_{j}}\}$ such that $f^{k_{j}}\to g$ in the compact-open topology. We denote the set of all such limit maps by $\Gamma(f)$. The target set of $f$, denoted by $T(f)$, is defined as the union of the images of all maps in $\Gamma(f)$:
$T(f):=\bigcup_{g\,\in\,\Gamma(f)}g(\Omega).$
Our first lemma will be useful in showing that the limit set of a holomorphic self-map — whose iterates are compactly divergent — is nonempty.
Lemma 4.5.
Let $\Omega\subsetneq\mathcal{X}$ be a Kobayashi hyperbolic domain in a complex manifold $\mathcal{X}$ that satisfies
$\liminf_{(z,\,w)\to(x,\,y)}\mathsf{k}_{\Omega}(z,\,w)>0\ \ \ \text{for every $x\neq y\in\partial\overline{\Omega}^{\rm End}$}.$ (4.2)
Let $f:\Omega\longrightarrow\Omega$ be a holomorphic map, and $\xi\in\partial\overline{\Omega}^{\rm End}$ be such that for every $z\in\Omega$, $f^{n}(z)\to\xi$. Then $f^{n}\to\xi$ in the compact-open topology.
Proof.
It is sufficient to show that given $p\in\Omega$, $r>0$ such that $B_{\mathsf{k}}(p,r)\subset\subset\Omega$, and a neighborhood $U_{\xi}$ of $\xi$, there exists $N_{0}\in\mathbb{N}$ such that
$f^{n}(B_{\mathsf{k}}(p,r))\subset U_{\xi}\ \ \ \text{for all $n\geq N_{0}$}.$
Case 1. $\xi\in\partial\overline{\Omega}^{\rm End}$ is an end.
Fix an exhaustion $(K_{j})_{j\geq 1}$ of $\overline{\Omega}$ by compact sets as described in Subsection 2.1. Suppose $\xi=(F_{j})_{j\geq 1}$ where each $F_{j}$ is a connected component of $\overline{\Omega}\setminus K_{j}$. Then we can find a $j_{0}\in\mathbb{N}$ such that $\widehat{F}_{j_{0}}\subset U_{\xi}$ where $(\widehat{F}_{j})_{j\geq 1}$ is the neighborhood basis of $\xi$ as described in Subsection 2.1. Clearly, there exists $N_{1}(p,j_{0})\in\mathbb{N}$ such that
$f^{n}(p)\in F_{j_{0}}\subset\widehat{F}_{j_{0}}\ \ \ \text{for all $n\geq N_{1}(p,j_{0})$}.$ (4.3)
Claim. There exists $N_{2}(p,j_{0})\in\mathbb{N}$ such that
$f^{n}(B_{\mathsf{k}}(p,r))\cap K_{j_{0}}=\emptyset\ \ \ \text{for all $n\geq N_{2}(p,j_{0})$}.$ (4.4)
Suppose that the above is not true, then there exist $(n_{j})_{j\geq 1}$ and $(x_{j})_{j\geq 1}\subset B_{\mathsf{k}}(p,r)$ such that $f^{n_{j}}(x_{j})\in K_{j_{0}}$. Extract a subsequence $(x_{j_{\nu}})$ such that $x_{j_{\nu}}\to x_{0}\in\overline{B_{\mathsf{k}}(p,r)}$. Then
$\mathsf{k}_{\Omega}(f^{n_{j_{\nu}}}(x_{j_{\nu}}),f^{n_{j_{\nu}}}(x_{0}))\leq\mathsf{k}_{\Omega}(x_{j_{\nu}},x_{0})\to 0,\ \ \ \text{as $\nu\to\infty$}.$ (4.5)
The limit points of the sequence $(f^{n_{j_{\nu}}}(x_{j_{\nu}}))_{\nu\geq 1}$ are in $\overline{\Omega}\cap K_{j_{0}}$. Since $f^{n_{j_{\nu}}}(x_{0})\to\xi$, (4.5) implies they are in $\partial\Omega$. But that contradicts our hypothesis (4.2). Hence the claim follows.
Let $N_{0}=\max\{N_{1}(p,j_{0}),N_{2}(p,j_{0})\}$ then for any $n\geq N_{0}$, both (4.3) and (4.4) are satisfied. Note that, for each $n\geq N_{0}$, $f^{n}(B_{\mathsf{k}}(p,r))$ is a connected set, disjoint from $K_{j_{0}}$ and intersecting $F_{j_{0}}$. This implies that
$f^{n}(B_{\mathsf{k}}(p,r))\subset F_{j_{0}}\subset U_{\xi}\ \ \ \text{for all $n\geq N_{0}$}.$
Hence we are done in this case.
Case 2. $\xi\in\partial\Omega$.
Choose a hermitian metric $h$ on $\mathcal{X}$ and let $R>0$ be small enough such that $B_{d_{h}}(\xi,R)\cap\overline{\Omega}\subset U_{\xi}$. Here, $B_{d_{h}}(\xi,R)$ denotes the open ball in $\mathcal{X}$, with respect to the distance induced by $h$, with center $\xi$ and radius $R>0$. We claim that there exists $N_{0}\in\mathbb{N}$ such that
$f^{n}(B_{\mathsf{k}}(p,r))\subset B_{d_{h}}(\xi,R)\ \ \ \text{for all $n\geq N_{0}$}.$
Suppose this is not true, then there exist $(n_{j})_{j\geq 1}$, and $(x_{j})\subset B_{\mathsf{k}}(p,r)$ such that $f^{n_{j}}(x_{j})\notin B_{d_{h}}(\xi,R)$ for all $j$. Passing to a subsequence, we can assume that $x_{j}\to x_{0}\in\overline{B_{\mathsf{k}}(p,r)}$. Therefore,
$\mathsf{k}_{\Omega}(f^{n_{j}}(x_{j}),f^{n_{j}}(x_{0}))\leq\mathsf{k}_{\Omega}(x_{j},x_{0})\to 0\ \ \ \text{as $j\to\infty$}.$
Since $f^{n_{j}}(x_{0})\to\xi\in\partial\Omega$, the limit points of the sequence $f^{n_{j}}(x_{j})$ must belong to $\partial\overline{\Omega}^{\rm End}\setminus B_{d_{h}}(\xi,R/2)$. But this will contradict (4.2) once again. Thus, our claim follows and we are done in this case too. ∎
The condition (4.2) in the above lemma is equivalent to $\Omega$ being hyperbolically embedded in $\mathcal{X}$. All domains that satisfy any type of visibility property satisfy this condition; see e.g. [13, Lemma 3.7]. Also, this condition is referred to as the boundary separation property in the latter article. We also need the following lemma.
Lemma 4.6.
Let $\Omega\subsetneq\mathcal{X}$ be a complete Kobayashi hyperbolic domain of a complex manifold $\mathcal{X}$. Suppose that
$\lim_{(z,w)\to(x,y)}\mathsf{k}_{\Omega}(z,w)=\infty\ \ \ \text{for any $x\neq y\in\partial\Omega$}.$ (4.6)
Then the above also holds for any two distinct $p,q\in\partial\overline{\Omega}^{\rm End}$.
Proof.
Given $p\neq q\in\partial\overline{\Omega}^{\rm End}$, if both are in $\partial\Omega$, we are trivially done. So assume that at least one of $p,q$ is an end. Now suppose (4.6) does not hold for $p,q$. Then there exist sequences $(z_{n})$, $(w_{n})$, $z_{n}\to p$ and $w_{n}\to q$ such that
$\sup\mathsf{k}_{\Omega}(z_{n},w_{n})=M<\infty.$
Let $\gamma_{n}:[a_{n},b_{n}]\longrightarrow\Omega$ be a Kobayashi geodesic joining $z_{n}$ and $w_{n}$. Then we claim that $(\gamma_{n})$ eventually avoids every compact $K\subset\Omega$. Otherwise, there exists a compact $K\subset\Omega$ such that — without loss of any generality — we may assume that each $\gamma_{n}$ intersects $K$. Let $t_{n}\in[a_{n},b_{n}]$ be such that $\gamma_{n}(t_{n})\in K$. Since $z_{n},\gamma_{n}(t_{n}),w_{n}\in{\rm range}(\gamma_{n})$ for each $n\in\mathbb{N}$, we have
$\mathsf{k}_{\Omega}(z_{n},K)\leq\mathsf{k}_{\Omega}(z_{n},\gamma_{n}(t_{n}))\leq\mathsf{k}_{\Omega}(z_{n},w_{n})<M,$
which implies that $(z_{n})$ lies in a bounded Kobayashi neighbourhood of the compact set $K$. The same holds for $(w_{n})$. Since $\Omega$ is complete hyperbolic, this implies that both $(z_{n})$, $(w_{n})$ lie in a compact subset of $\Omega$, which contradicts that $p,q\in\partial\overline{\Omega}^{\rm End}$.
Using the fact that $(\gamma_{n})$ eventually avoids every compact subset of $\Omega$, one can find points $x\neq y\in\partial\Omega$ and sequences $(s_{n})_{n\geq 1},(t_{n})_{n\geq 1}\subset[a_{n},b_{n}]$ such that $\gamma_{n}(s_{n})\to x$ and $\gamma_{n}(t_{n})\to y$. We refer the reader to Lemma 2.2 and the proof of Theorem 1.6 in [13] for the precise argument. Observe that
[…]
which contradicts our hypothesis, whence the result follows. ∎
4.3. The proofs of Theorem 1.5, Theorem 1.6 and Propostion 1.7.
We now have all the ingredients to present:
The proof of Theorem 1.5.
Fix a point $p\in\Omega$. Since $(f^{n})_{n\geq 1}$ is compactly divergent, the completeness of $\mathsf{k}_{\Omega}$ implies that
$\lim_{n\to\infty}{\mathsf{k}_{\Omega}\big{(}p,f^{n}(p)\big{)}}=\infty.$
There exists a strictly increasing sequence of natural numbers $(n_{\nu})_{\nu\geq 1}$ such that
$\mathsf{k}_{\Omega}(p,f^{n_{\nu}+1}(p))>\mathsf{k}_{\Omega}(p,f^{n_{\nu}}(p))\ \ \ \forall\nu\in\mathbb{N}.$ (4.7)
Let us write $w_{n_{\nu}}=f^{n_{\nu}}(p)$. Since $\overline{\Omega}^{\rm End}$ is sequentially compact, by passing to subsequences if necessary, we can assume that $w_{n_{\nu}}\to x\in\partial\overline{\Omega}^{\rm End}$ in the topology of end-point compactification, and $i_{\mathsf{k}}(w_{n_{\nu}})\to\xi\in\partial\overline{\Omega}^{\mathsf{k}}$ in the topology of metric compactification, as $\nu\to\infty$. Clearly, by definition, $\xi\in\mathscr{H}(x)$.
Claim 1: $f(w_{n_{\nu}})\to x$ as $\nu\to\infty$.
Suppose, to get a contradiction, that the above is not true. Then, up to passing to a subsequence, we can assume that $f(w_{n_{\nu}})\to y$ as $\nu\to\infty$ for some $x\neq y\in\partial\overline{\Omega}^{\rm End}$. Observe that
$\mathsf{k}_{\Omega}\big{(}w_{n_{\nu}},f(w_{n_{\nu}})\big{)}=\mathsf{k}_{\Omega}\big{(}f^{n_{\nu}}(p),f^{n_{\nu}+1}(p)\big{)}\leq\mathsf{k}_{\Omega}\big{(}p,f(p)\big{)}<\infty.$
Hence, $\liminf_{(z,w)\to(x,y)}{\mathsf{k}_{\Omega}(z,w)<\infty}$. However, Lemma 4.6 implies that this is not possible, whence the claim.
Once again, passing to a subsequence if necessary, we can further assume that $i_{\mathsf{k}}(f(w_{n_{\nu}}))\to\xi_{1}$ in the topology of metric compactification, where $\xi_{1}\in\mathscr{H}(x)$.
Claim 2: For any $R>0$, we have
$f(H_{p}(\xi,R)\subset H_{p}(\xi_{1},R).$ (4.8)
Recall, $H_{p}(\xi,R)$, $H_{p}(\xi_{1},R)$ are the horoballs centered at $\xi$, $\xi_{1}$ respectively, with respect to the pole $p$ and radius $R>0$. To see the above, fix an $R>0$ and let $z\in H_{p}(\xi,R)$. Then
$\mathsf{k}_{\Omega}\big{(}f(z),f(w_{n_{\nu}})\big{)}-\mathsf{k}_{\Omega}\big{(}p,f(w_{n_{\nu}})\big{)}$ $\leq\mathsf{k}_{\Omega}(z,w_{n_{\nu}})-\mathsf{k}_{\Omega}(p,f(w_{n_{\nu}}))$
$=\mathsf{k}_{\Omega}(z,w_{n_{\nu}})-\mathsf{k}_{\Omega}(p,w_{n_{\nu}})+\mathsf{k}_{\Omega}(p,w_{n_{\nu}})-\mathsf{k}_{\Omega}(p,f(w_{n_{\nu}}))$
$\leq\mathsf{k}_{\Omega}(z,w_{n_{\nu}})-\mathsf{k}_{\Omega}(p,w_{n_{\nu}}),$ (4.9)
where the last inequality above follows from (4.7). It follows from (4.3) that
$h_{p,\,\xi_{1}}(f(z))\leq h_{p,\,\xi}(z)<\tfrac{1}{2}\log R,$
where $h_{p,\,\xi_{1}}$ and $h_{p,\,\xi}$ are the unique horofunctions representing $\xi_{1},\xi$ respectively, and that vanish at the point $p$. From this our claim follows.
Applying the method of the proof of Claim 1 and Claim 2 above, inductively, we see that for each $n\in\mathbb{N}$, there exists $\xi_{n}\in\mathscr{H}(x)$ such that
$f^{n}(H_{p}(\xi,R))\subset H_{p}(\xi_{n},R)\ \ \ \text{for all $R>0$}.$
Since $H^{b}_{p}(x,R)$ is the union of all the horoballs $H_{p}(\eta,R)$, $\eta\in\mathscr{H}(x)$, the above inclusion implies (1.2). Note that $H^{s}_{p}(x,R)$ is the intersection of all horoballs $H_{p}(\eta,R)$, $\eta\in\mathscr{H}(x)$, and hence is contained in $H_{p}(\xi,R)$, from which the inclusion (1.3) follows. ∎
As mentioned in Section 1, after the statement of Theorem 1.5, the property that iterates map a certain horoball into the big horosphere associated to it, helps us in proving the following result about the location of the target set.
Lemma 4.7.
Let $\Omega\subsetneq\mathcal{X}$ be a complete Kobayashi hyperbolic domain of a complex manifold $\mathcal{X}$ satisfying property (4.6). Let $f:\Omega\longrightarrow\Omega$ be a holomorphic map such that $(f^{n})_{n\geq 1}$ is compactly divergent. Then $\Gamma(f)$ is nonempty and each $h\in\Gamma(f)$ is a constant function. Moreover,
$T(f)\subset\bigcap_{R>0}\,\overline{H^{b}_{p}(x,R)},$
where $p\in\Omega$ is a fixed point, and $x\in\partial\overline{\Omega}^{\rm End}$ is as given by Theorem 1.5.
Proof.
Since $(f^{n})_{n\geq 1}$ is compactly divergent and $(\Omega,\mathsf{k}_{\Omega})$ is complete, given $p\in\Omega$, we have
$\lim_{n\to\infty}\mathsf{k}_{\Omega}(p,f^{n}(p))=\infty.$
As $\overline{\Omega}^{\rm End}$ is sequentially compact, there exist a a subsequence $(f^{n_{j}})_{j\geq 1}$ and a point $y\in\partial\overline{\Omega}^{\rm End}$ such that $f^{n_{j}}(p)\to y$. We claim that $f^{n_{j}}(z)\to y$ for every $z\in\Omega$. Consider a limit point $w\in\partial\overline{\Omega}^{\rm End}$ of the sequence $(f^{n_{j}}(z))$. Without loss of generality, we may assume that $f^{n_{j}}(z)\to w$. Since
$\mathsf{k}_{\Omega}(f^{n_{j}}(z),f^{n_{j}}(p))\leq\mathsf{k}_{\Omega}(z,p)<\infty,$
it follows from our hypothesis that $w=y$. This implies that the only limit point of $(f^{n_{j}}(z))$ is $y$. Now an application of Lemma 4.5 implies that $f^{n_{j}}\to y$ in the compact-open topology, and hence, $y\in\Gamma(f)$. Here, $y$ also denotes the constant map.
Let $g\in\Gamma(f)$, clearly $g(\Omega)\subset\partial\overline{\Omega}^{\rm End}$. Suppose $g$ is not constant, and let $y_{1}\neq y_{2}\in\partial\overline{\Omega}^{\rm End}$ be such that $y_{1}=g(z)$, $y_{2}=g(w)$. Let $(f^{k_{j}})$ be such that $f^{k_{j}}\to g$ in the compact-open topology. Then
$\mathsf{k}_{\Omega}(f^{k_{j}}(z),f^{k_{j}}(w))\leq\mathsf{k}_{\Omega}(z,w)\ \ \ \forall j\in\mathbb{N}.$
Since $f^{k_{j}}(z)\to y_{1}$ and $f^{k_{j}}(w)\to y_{2}$, the above inequality contradicts our hypothesis. Therefore, $g$ is constant.
Let $x\in\partial\overline{\Omega}^{\rm End}$ and $\xi\in\mathscr{H}(x)$ be as given by Theorem 1.5. Then for any $R>0$, and any $n\in\mathbb{N}$, we have
$f^{n}(H_{p}(\xi,R))\subset H^{b}_{p}(x,R)\ \ \ \forall R>0.$
Observe that given $g\in\Gamma(f)$, since $g$ is a constant function, $g(\Omega)=g(H_{p}(\xi,R))$ for any $R>0$. Therefore, for each $R>0$ we have
$g(\Omega)=g(H_{p}(\xi,R))=\lim_{j\to\infty}f^{k_{j}}(H_{p}(\xi,R))\subset\overline{H^{b}_{p}(x,R)}\cap\partial\overline{\Omega}^{\rm End}.$
This implies
$g(\Omega)\subset\bigcap_{R>0}\big{(}\overline{H^{b}_{p}(x,R)}\cap\partial\overline{\Omega}^{\rm End}\big{)},$
and since $g\in\Gamma(f)$ is arbitrary, the result follows. ∎
We are now ready to present the proof of Theorem 1.6.
The proof of Theorem 1.6.
[…]
$x_{g}\in\bigcap_{R>0}\overline{H^{b}_{p}(x,R)}$
for some $x\in\partial\overline{\Omega}^{\rm End}$. By our assumption, there exists a $y\in\partial\overline{\Omega}^{\rm End}$ such that
$\bigcap_{R>0}\overline{H^{b}_{p}(x,R)}=\{y\},$
therefore, $g\equiv y$, where $g\in\Gamma(f)$ is arbitrary. From this the result follows. ∎
We now present the proof of Proposition 1.7 but first we recall the following result.
Result 4.8 (paraphrasing [3, Lemma 8]).
Let $D\Subset\mathbb{C}^{d}$ be a bounded convex domain. Let $f:D\longrightarrow D$ be a holomorphic map without a fixed point. Then there exist $x\in\partial D$ and $\xi\in\mathscr{H}(x)$ such that, for a fixed point $p\in D$, we have
$f(H_{p}(\xi,R))\subset H_{p}(\xi,R)\quad\forall R>0.$
Abate–Raissy stated the above result in terms of a horosphere sequence $(x_{\nu})$ converging to $x$. In our language, this basically means $i_{\mathsf{k}}(x_{\nu})\to\xi$ for some $\xi\in\mathscr{H}(x)$. We are now ready to present:
The proof of Proposition 1.7.
Suppose $f$ does not have any fixed point. Then by [1, Theorem 2.4.20] the sequence $(f^{n})$ is compactly divergent. Appealing to Montel’s Theorem, as $\Omega$ is bounded, we see that $\Gamma(f)$ is nonempty. Moreover, the boundary divergence condition 4.6 implies — as shown in the proof of Lemma 4.7 — that each function in $\Gamma(f)$ is constant. Now regarding the target set, we first observe that by Result 4.8, there exist $x\in\partial D$ and $\xi\in\mathscr{H}(x)$ such that, for a fixed point $p\in D$, and for any $n\in\mathbb{N}$, we have
$f^{n}(H_{p}(\xi,R))\subset H_{p}(\xi,R)\quad\forall R>0.$
Exactly proceeding as in the proof of Lemma 4.7, we see that the target set
$T(f)\subset\bigcap_{R>0}\overline{H_{p}(\xi,\,R)}.$
By our assumption, the set in the right-side of the above inclusion is $\{x\}$. (Observe that Proposition 4.2 implies that this set contains $x$). This completes the proof. ∎
5. Continuous extension of the canonical map
In this section, we present the proof of Theorem 1.8. We also establish a necessary condition and a sufficient condition for the existence of the continuous map $F:\overline{\Omega}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\rm End}$ such that
$F\circ i_{\mathsf{k}}={i}_{\Omega}\ \ \ \text{on $\Omega$}.$ (5.1)
In particular, we observe that the property ensuring such an extension appears to be weaker than the geodesic visibility property.
5.1. The proof of Theorem 1.8
Proof.
Suppose there exists a continuous surjective map $F:\overline{\Omega}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\rm End}$ that satisfies (5.1).
Claim. For any $x\in\partial\overline{\Omega}^{\rm End}$, $F^{-1}(x)=\mathscr{H}(x)$.
Let $\xi\in F^{-1}(x)$ and let $(x_{n})\subset\Omega$ be such that $i_{\mathsf{k}}(x_{n})\to\xi$. This implies that $F\circ i_{\mathsf{k}}(x_{n})\to F(\xi)$. Since $F$ satisfies (5.1), and $F(\xi)=x$, we get that $x_{n}\to x$. Since $(x_{n})$ has no limit point in $\Omega$, it follows — by Remark 3.2 — that $\xi\in\partial\overline{\Omega}^{\mathsf{k}}$ whence, by definition, $\xi\in\mathscr{H}(x)$ and $F^{-1}(x)\subset\mathscr{H}(x)$. Conversely, suppose $\xi\in\mathscr{H}(x)$ then there exists a sequence $(x_{n})\subset\Omega$ such that $x_{n}\to x$ and $i_{\mathsf{k}}(x_{n})\to\xi$. Therefore, $F\circ i_{\mathsf{k}}(x_{n})\to F(\xi)$ but $F\circ i_{\mathsf{k}}(x_{n})=x_{n}$, and since $x_{n}\to x$, we get that $F(\xi)=x$. So $F(\mathscr{H}(x))\subset\{x\}$ for all $x\in\partial\overline{\Omega}^{\rm End}$. This establishes the claim from which the only if part of the statement of the theorem follows.
Suppose now $\mathscr{H}(x)\cap\mathscr{H}(y)=\emptyset$ for all $x\neq y\in\partial\overline{\Omega}^{\rm End}$. Given $\xi\in\partial\overline{\Omega}^{\mathsf{k}}$, there exists a sequence $(x_{n})_{n\geq 1}$ such that $i_{\mathsf{k}}(x_{n})\to\xi$ in the topology of $\overline{\Omega}^{\mathsf{k}}$. Note that — by Remark 3.2 — $(x_{n})_{n\geq 1}$ does not have a limit point in $\Omega$. Since $\overline{\Omega}^{\rm End}$ is sequentially compact, there exists a sequence $(x_{n_{k}})_{k\geq 1}$ such that $x_{n_{k}}\to x\in\partial\overline{\Omega}^{\rm End}$, as $k\to\infty$, in the topology of $\overline{\Omega}^{\rm End}$. This implies, by definition, $\xi\in\mathscr{H}(x)$, i.e., we have
$\partial\overline{\Omega}^{\mathsf{k}}=\overline{\Omega}^{\mathsf{k}}\setminus i_{\mathsf{k}}(\Omega)=\bigsqcup_{x\in\partial\overline{\Omega}^{\rm End}}\mathscr{H}(x).$ (5.2)
Consider the map $F:\overline{\Omega}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\rm End}$ defined by
$F(\xi):=\begin{cases}i_{\mathsf{k}}^{-1}(\xi),&\ \ \ \xi\in i_{\mathsf{k}}(\Omega),\\ x,&\ \ \ \text{if $\xi\in\mathscr{H}(x)$.}\end{cases}$
Note $F$ is a well-defined, surjective map that satisfies condition (5.1). We now show that $F$ is continuous. It is sufficient to show that $F$ is continuous on $\partial\overline{\Omega}^{\mathsf{k}}$. Let $(\xi_{n})_{n\geq 1}\subset\overline{\Omega}^{\mathsf{k}}$ be such that $\xi_{n}\to\xi\in\partial\overline{\Omega}^{\mathsf{k}}$, in the topology of $\overline{\Omega}^{\mathsf{k}}$. We may assume, without loss of any generality, that either $(\xi_{n})_{n\geq 1}\subset i_{\mathsf{k}}(\Omega)$ or $(\xi_{n})_{n\geq 1}\subset\partial\overline{\Omega}^{\mathsf{k}}$.
Case 1. $(\xi_{n})_{n\geq 1}\subset i_{\mathsf{k}}(\Omega)$.
In this case, let $(x_{n})_{n\geq 1}\subset\Omega$ be such that $x_{n}=F(\xi_{n})=i_{\mathsf{k}}^{-1}(\xi_{n})$ for all $n$. Also, let $F(\xi)=x\in\partial\overline{\Omega}^{\rm End}$, i.e., there exists a sequence $(z_{n})_{n\geq 1}\subset\Omega$ such that $z_{n}\to x$ and $i_{\mathsf{k}}(z_{n})\to\xi$. Let $y$ be a limit point of $(x_{n})_{n\geq 1}$, since $i_{\mathsf{k}}(x_{n})=\xi_{n}\to\xi\in\partial\overline{\Omega}^{\mathsf{k}}$, $y\in\partial\overline{\Omega}^{\rm End}$. Let $(x_{n_{k}})_{k\geq 1}$ be such that $x_{n_{k}}\to y\in\partial\overline{\Omega}^{\rm End}$. Note $i_{\mathsf{k}}(x_{n_{k}})\to\xi$ and hence by definition $\xi\in\mathscr{H}(y)\cap\mathscr{H}(x)$. By our assumption, we have $y=x$. Since $y$ is arbitrary, $(x_{n})_{n\geq 1}$ converges to $x$ and therefore $F(\xi_{n})$ converges to $F(\xi)$, and hence we are done.
Case 2. $(\xi_{n})_{n\geq 1}\subset\partial\overline{\Omega}^{\mathsf{k}}$.
Let $F(\xi_{n})=e_{n}\in\partial\overline{\Omega}^{\rm End}$. To show that $F(\xi_{n})$ converges to $F(\xi)$, we will show that if $e\in\partial\overline{\Omega}^{\rm End}$ is a limit point of $(e_{n})$, then $e=F(\xi)$. Without loss of any generality, we may assume that $e_{n}\to e\in\partial\overline{\Omega}^{\rm End}$. Note that, by the definition of $F$, for each $n$, there exists a sequence $(z^{n}_{\nu})_{\nu\geq 1}$ such that $z^{n}_{\nu}\to e_{n}$ and $i_{\mathsf{k}}(z^{n}_{\nu})\to\xi_{n}$ as $\nu\to\infty$.
Subcase (a). $e\in\partial\Omega$, i.e., $e$ is not an end.
Since $e_{n}\to e$ then except finitely many $n$, all $e_{n}\in\partial\Omega$. Without loss of any generality, assume $e_{n}\in\partial\Omega$ for all $n$ and $e_{n}\to e$, as $n\to\infty$. Fix $n\in\mathbb{N}$, since $z^{n}_{\nu}\to e_{n}$ as $\nu\to\infty$, we can find $N(n)\in\mathbb{N}$ such that $d_{h}(z^{n}_{\nu},e_{n})<1/n$ for all $\nu\geq N(n)$. Also, since $i_{\mathsf{k}}(z^{n}_{\nu})\to\xi_{n}$, as $\nu\to\infty$, we can find $M(n)\geq N(n)$ such that $d_{\mathsf{k}}(\xi_{n},i_{\mathsf{k}}(z^{n}_{\nu}))<1/n$ for all $\nu\geq M(n)$. Here, $d_{\mathsf{k}}$ is any distance that metrizes the space $\overline{\Omega}^{\mathsf{k}}$. Consider the sequence given by $x_{n}=z^{n}_{M(n)}$ and observe since
$d_{h}(x_{n},e)$ $\leq d_{h}(x_{n},e_{n})+d_{h}(e_{n},e)<1/n+d_{h}(e_{n},e)\ \ \ \text{and}$
$d_{\mathsf{k}}(i_{\mathsf{k}}(x_{n}),\xi)$ $\leq d_{\mathsf{k}}(i_{\mathsf{k}}(x_{n}),\xi_{n})+d_{\mathsf{k}}(\xi_{n},\xi)<1/n+d_{\mathsf{k}}(\xi_{n},\xi).$
It follows that $x_{n}\to e$ and $i_{\mathsf{k}}(x_{n})\to\xi$ as $n\to\infty$, whence $F(\xi)=e$.
Subcase (b) $e$ is an end, and $e_{n}\in\partial\Omega$ for all $n\in\mathbb{N}$.
We choose and fix an exhaustion $(K_{j})_{j\geq 1}$ of $\overline{\Omega}$ by compact sets. Let $e=(F_{j})_{j\geq 1}$ where, for each $j$, $F_{j}$ is a connected component of $\overline{\Omega}\setminus K_{j}$ such that $F_{j+1}\subset F_{j}$. Recall the neighborhood basis for $e$ is given by the family $(\widehat{F}_{j})_{j\geq 1}$ of subsets of $\overline{\Omega}^{\rm End}$ where
$\widehat{F}_{j}:=F_{j}\cup\{f\,|\,\text{$f=(G_{\nu})_{\nu\geq 1}$ is an end of $\overline{\Omega}$ such that $G_{\nu}=F_{\nu}\ \ \forall\nu=1,\ldots,j$}\}.$ (5.3)
Fix $j$, since $\widehat{F}_{j}$ is a neighborhood of $e$, there exists $N(j)\in\mathbb{N}$ such that $e_{n}\in\widehat{F}_{j}$ for all $n\geq N(j)$. Since $e_{n}\in\partial\Omega$, we have $e_{n}\in F_{j}$ for all $n\geq N(j)$. By Lemma 2.1, for each $j$, $F_{j}$ is a neighborhood of $e_{n}$ for all $n\geq N(j)$. Therefore, for each $j$, we can find large enough $M(j)\in\mathbb{N}$ such that
$x_{j}:=z^{j+N(j)}_{M(j)}\in F_{j}\ \ \ \text{and \ \ $d_{\mathsf{k}}\Big{(}\xi_{j+N(j)},i_{\mathsf{k}}\big{(}z^{j+N(j)}_{M(j)}\big{)}\Big{)}<\frac{1}{j}$}.$
Clearly $x_{j}\to e$ and $i_{\mathsf{k}}(x_{j})\to\xi$ as $j\to\infty$. Therefore, in this case too we have $F(\xi)=e$.
Subcase (c). $e$ and $e_{n}$’s, $n\in\mathbb{N}$, are ends.
Fix $n\in\mathbb{N}$, since $i_{\mathsf{k}}(z^{n}_{\nu})\to\xi_{n}$ as $\nu\to\infty$, we can find $N_{1}(n)\in\mathbb{N}$ such that
$d_{\mathsf{k}}\big{(}i_{\mathsf{k}}(z^{n}_{\nu}),\xi_{n}\big{)}<\frac{1}{n}\ \ \ \forall\nu\geq N_{1}(n).$
Moreover, we can also assume that $N_{1}(n)$ is strictly increasing as a function of $n$. Since $e_{n}$’s are end, we write $e_{n}=(E^{n}_{\nu})_{\nu\geq 1}$. For each $n$, since $\widehat{E}^{n}_{n}$ is a neighborhood of $e_{n}$ and $z^{n}_{\nu}\to e_{n}$, we can find $N_{2}(n)$, again strictly incresing in $n$, such that $z^{n}_{\nu}\in\widehat{E}^{n}_{n}$ for all $\nu\geq N_{2}(n)$. Let $N(n):=\max\{N_{1}(n),N_{2}(n)\}$ and consider the sequence $x_{n}=z^{n}_{N(n)}$. Clearly $i_{\mathsf{k}}(x_{n})\to\xi$ and $x_{n}\in E^{n}_{n}$ for all $n$. We now show that $x_{n}\to e$ as $n\to\infty$. Recall the neighborhood basis for $e$ is given by the family $(\widehat{F}_{j})_{j\geq 1}$ of subsets of $\overline{\Omega}^{\rm End}$ as in (5.3).
Claim: For each $j\geq 1$, there exists $N_{j}\in\mathbb{N}$ such that $x_{n}\in\widehat{F}_{j}$ for all $n\geq N_{j}$.
Fix $j\geq 1$, since $e_{n}\to e$ as $n\to\infty$, there exists a $N_{j}$ such that $e_{n}\in\widehat{F}_{j}$ for all $n\geq N_{j}\geq j$. Choose $n\geq N_{j}$ arbitrary then $e_{n}\in\widehat{F}_{j}$ implies that the end $e_{n}=(E^{n}_{\nu})_{\nu\geq 1}$ satisfies $E^{n}_{\nu}=F_{\nu}$ for $\nu=1,2,...,j$. For $n\geq N_{j}\geq j$, we have $x_{n}\in E^{n}_{n}\subset E^{n}_{j}=F_{j}\subset\widehat{F}_{j}$. Hence the claim follows.
This implies that $x_{n}\to e$ as $n\to\infty$, and therefore, $F(\xi)=e$ in this case too. This completes the proof. ∎
In complete hyperbolic domains, a necessary condition for the existence of the map $F$ satisfying condition 5.1, as in the above theorem, is given by the following proposition.
Proposition 5.1.
Let $\Omega\subsetneq\mathcal{X}$ be a complete Kobayashi hyperbolic domain in a complex manifold $\mathcal{X}$. Suppose there exists a continuous surjective map $F:\overline{\Omega}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\mathrm{End}}$ such that
$F\circ i_{\mathsf{k}}=i_{\Omega}\quad\text{on }\Omega.$
Then every geodesic ray $\gamma:\mathbb{R}_{+}\longrightarrow\Omega$ lands in $\overline{\Omega}^{\mathrm{End}}$.
Proof.
Given a geodesic ray $\gamma$, consider the family of functions
$\left(\mathsf{k}_{\Omega}(\cdot,\gamma(t))-\mathsf{k}_{\Omega}(\gamma(t),\gamma(0))\right)_{t\geq 0}.$
Each function in the family is 1-Lipschitz and the family is monotonically decreasing (in $t$) and locally bounded above. Hence, by the Arzelà–Ascoli theorem, the family converges uniformly on compact subsets to a 1-Lipschitz function $F_{\gamma}$, given by
$F_{\gamma}(\cdot):=\lim_{t\to\infty}\left(\mathsf{k}_{\Omega}(\cdot,\gamma(t))-\mathsf{k}_{\Omega}(\gamma(t),\gamma(0))\right).$
This function defines a horofunction, and let $\xi:=[F_{\gamma}]$ denote its equivalence class. Note that in our notation, $h_{p,\,\xi}(\cdot)=F_{\gamma}(\cdot)$, where $p=\gamma(0)$. Define the cluster set of $\gamma$ in $\overline{\Omega}^{\mathrm{End}}$ as
$C(\gamma):=\left\{x\in\partial\overline{\Omega}^{\mathrm{End}}\;\middle|\;\exists\,(t_{n})\to\infty\text{ such that }\gamma(t_{n})\to x\right\}.$
For any $x\in C(\gamma)$, we must have $\xi\in\mathscr{H}(x)$ by the construction. Hence,
$\xi\in\bigcap_{x\in C(\gamma)}\mathscr{H}(x).$
By Theorem 1.8, we know that $\mathscr{H}(x)\cap\mathscr{H}(y)=\emptyset$ whenever $x\neq y\in\overline{\Omega}^{\mathrm{End}}$. It follows that $C(\gamma)$ must be a singleton. Therefore, the geodesic ray $\gamma$ lands at a unique point in $\partial\overline{\Omega}^{\mathrm{End}}$. ∎
We remind the reader that the geodesic visibility property also ensures that geodesic rays land at a point in the boundary, see [8, Lemma 3.1]. However, our next result — which provides a sufficient condition for the existence of the map $F$ in Theorem 1.8 — shows that geodesic visibility alone suffices to guarantee such an extension.
Proposition 5.2.
Let $\Omega\subsetneq\mathcal{X}$ be a complete Kobayashi hyperbolic domain. Suppose for a fixed $p\in\Omega$, and any $x\in\partial\overline{\Omega}^{\rm End}$ we have
$\bigcap_{R>0}\overline{H^{b}_{p}(x,R)}=\{x\}.$ (5.4)
Then, for any $\xi\in\mathscr{H}(x)$, we also have
$\bigcap_{R>0}\overline{H_{p}(\xi,R)}=\{x\}.$ (5.5)
In particular, there exists a continuous surjective map $F:\overline{\Omega}^{\mathsf{k}}\longrightarrow\overline{\Omega}^{\rm End}$ such that
$F\circ i_{\mathsf{k}}={i}_{\Omega}\ \ \ \text{on \ \ $\Omega$}.$
Proof.
Given $x\in\partial\overline{\Omega}^{\rm End}$, for any $R>0$, we have
$H^{b}_{p}(x,R)=\bigcup_{\xi\in\mathscr{H}(x)}H_{p}(\xi,R),$
where, we recall: $H_{p}(\xi,R):=\big{\{}z\in\Omega:h_{p,\,\xi}(z)<\tfrac{1}{2}\log R\big{\}}$. Here, $h_{p,\,\xi}$ is the unique horofunction representing $\xi$ and vanishing at $p$. Therefore, for any $R>0$,
$\overline{H_{p}(\xi,R)}\subset\overline{H^{b}_{p}(x,R)}\ \ \ \text{$\implies$}\ \ \ \bigcap_{R>0}\overline{H_{p}(\xi,R)}\subset\bigcap_{R>0}\overline{H^{b}_{p}(x,R)}=\{x\}$ (5.6)
We now claim that $\bigcap_{R>0}\overline{H_{p}(\xi,R)}$ is non-empty. To see that, for each $n$, choose $z_{n}\in H_{p}(\xi,1/n)$. This is possible since for any $R>0$, each $H_{p}(\xi,R)$ is non-empty, according to Lemma 3.4. By passing to a subsequence, we can assume $z_{n}\to p\in\partial\overline{\Omega}^{\rm End}$. Note that for any $R>0$, there exists $N(R)$ such that $z_{n}\in\overline{H_{p}(\xi,R)}$ for all $n\geq N(R)$. Therefore, $p\in\overline{H_{p}(\xi,R)}$ for all $R>0$ whence the claim.
Therefore, the inclusion (5.6) and the above claim implies (5.5). Let $x\neq y\in\partial\overline{\Omega}^{\rm End}$ and let $\xi\in\mathscr{H}(x)\cap\mathscr{H}(y)$. Then, it follows, from (5.5) that
$\bigcap_{R>0}\overline{H_{p}(\xi,R)}=\{x\}=\{y\},$
which is a contradiction. Hence, $\mathscr{H}(x)\cap\mathscr{H}(y)=\emptyset$ whenever $x\neq y\in\partial\overline{\Omega}^{\rm End}$. The result now follows from an application of Theorem 1.8. ∎
As noted earlier, both geodesic visibility domains and strictly convex domains satisfy condition (5.4). Therefore, it follows from the above proposition that the canonical map extends as a continuous map from $\overline{\Omega}^{\mathsf{k}}$ to $\overline{\Omega}^{\rm End}$. Note that, it also implies that, geodesic rays land in a strictly convex domains.
We also record the following result, a part of which — namely, the equivalence of (1) and (3) — is known in the setting of bounded domains in $\mathbb{C}^{n}$. For completeness, we briefly sketch a proof in our general setting.
Theorem 5.3.
Let $\Omega\subsetneq\mathcal{X}$ be a Kobayashi hyperbolic domain of a complex manifold $\mathcal{X}$ and let $o\in\Omega$ be a fixed point. Then the following are equivalent:
  1. (1)
    For every $x\in\partial\overline{\Omega}^{\rm End}$, the limit
    $\lim_{w\to x}(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w))\ \ \ \text{exists for all $z\in\Omega$}.$
  2. (2)
    For every $x\in\partial\overline{\Omega}^{\rm End}$, the set $\mathscr{H}(x)$ is a singleton set.
  3. (3)
    The map $i_{\mathsf{k}}$ extends continuously from $\overline{\Omega}^{\rm End}\longrightarrow\overline{\Omega}^{\mathsf{k}}$.
Proof.
Suppose (1)(1)( 1 ) holds. Fix $x\in\partial\overline{\Omega}^{\rm End}$ and let $\xi_{1},\xi_{2}\in\mathscr{H}(x)$. Then there exist sequences $(x_{n}),(y_{n})$, both converging to $x$, such that $[h_{o,\,\xi_{1}}]=\xi_{1}$ and $[h_{o,\,\xi_{2}}]=\xi_{2}$, where
$h_{o,\,\xi_{1}}(\cdot)$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(o,x_{n})\big{)},$
$h_{o,\,\xi_{2}}(\cdot)$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,y_{n})-\mathsf{k}_{\Omega}(o,y_{n})\big{)}$
are the horofunctions representing $\xi_{1},\xi_{2}$, and which vanish at $o$. Clearly, by our assumption $h_{o,\,\xi_{1}}(\cdot)=h_{o,\,\xi_{2}}(\cdot)$ and hence $\xi_{1}=\xi_{2}$. This implies $\mathscr{H}(x)$ is singleton and since $x\in\partial\overline{\Omega}^{\rm End}$ is arbitrary, we have (2)(2)( 2 ). Now suppose (2)(2)( 2 ) holds for each $x\in\partial\overline{\Omega}^{\rm End}$. Fix an $x\in\partial\overline{\Omega}^{\rm End}$ and suppose there exists $z\in\Omega$ such that
$\liminf_{w\to x}(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w))\neq\limsup_{w\to x}(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w))$ (5.7)
Let $(x_{n}),(y_{n})$ be sequences, both converging to $x$, such that
$\liminf_{w\to x}(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w))$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(z,x_{n})-\mathsf{k}_{\Omega}(o,x_{n})\big{)}$
$\limsup_{w\to x}(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w))$ $=\lim_{n\to\infty}(\mathsf{k}_{\Omega}(z,y_{n})-\mathsf{k}_{\Omega}(o,y_{n})).$
By passing to subsequences, we can find $\xi_{1},\xi_{2}\in\mathscr{H}(x)$ such that
$h_{o,\,\xi_{1}}(\cdot)$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,x_{n})-\mathsf{k}_{\Omega}(o,x_{n})\big{)},$
$h_{o,\,\xi_{2}}(\cdot)$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\cdot,y_{n})-\mathsf{k}_{\Omega}(o,y_{n})\big{)}.$
Note that $\xi_{1}=\xi_{2}$ if and only if $h_{o,\,\xi_{1}}(\cdot)=h_{o,\,\xi_{2}}(\cdot)$ which is not the case due to (5.7). Therefore, $\xi_{1}\neq\xi_{2}\in\mathscr{H}(x)$, a contradiction to our assumtion. Hence (2)(2)( 2 ) implies (1)(1)( 1 ).
Observe that (3)(3)( 3 ) clearly implies (2)(2)( 2 ). Now suppose (2)(2)( 2 ) holds and consider the map $G:\overline{\Omega}^{\rm End}\longrightarrow\overline{\Omega}^{\mathsf{k}}$ defined by
$G(x):=\begin{cases}i_{\mathsf{k}}(x),&\ \ \ x\in\Omega,\\ \xi,&\ \ \ \text{otherwise, where $\mathscr{H}(x)=\{\xi\}$.}\end{cases}$
It is a routine — following the same strategy as in the proof of Theorem 1.8 — to check that $G$ is a continuous map. ∎